OpenStudy (anonymous):

how do you rewrite an equation for using substitution method that already isnt in y= form? Please give example step by step

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

Are you talking about a system of equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I believe it is solve each equation for a variable and substitute on in the other.

OpenStudy (anonymous):

yes! like how would you solve.. 2x – 3y = –2 4x + y = 24 the top equation getting y by itself so I can substitute?

OpenStudy (anonymous):

Better to solve the bottom equation for y and substitute it into the top equation. The bottom equation can be solved for y just by moving the 4x to the other side, giving you y=24 -4x. Then just take that and substitute it into the top equation for y.

OpenStudy (anonymous):

x + y = 11 3x - y = 5 so here's a sample system of equations. first you have to solve for one variable, like evansda said. so let's solve for y. x + y = 11 y = 11 - x Then we substitute in this y value into the second equation. 3x - (11 - x) = 5 Now solve. 3x - (11 - x) = 5 3x - 11 + x = 5 4x - 11 = 5 4x = 16 x = 4 Now insert this x value into either equation and solve for y. x + y = 11 4 + y = 11 y = 7 So we get x=4, y = 7 or (4,7)

OpenStudy (anonymous):

thank you guys so much